OpenStudy (anonymous):

Solve for x by completing the square: 4x^2 + 6x - 6 = 0 A. (-3+√33)/4, (-3-√33)/4 B. (3+√33)/4, (3-√33)/4 C. 3, -2 D. 1, 6

OpenStudy (anonymous):

C

OpenStudy (wolf1728):

To complete the square: Move the "non X" term to the right: 4x² + 6x = 6 Divide the equation by the coefficient of X² x² + 1.5x = 1.5 take the coefficient of X which is 1.5 divide it by 2 making it .75 square that number .75² = .5625 then add it to both sides of the equation x² + 1.5x + .5625 = 1.5 + .5625 Take the square root of both sides (X + .75) * (x+.75) = sqrt (2.0625) (X + .75) * (x+.75) = 1.4361406616 x = -.75 -1.4361406616 x= -2.1861406616 Second Root x = -.75 + 1.4361406616 x = 0.6861406616 and the answer is A

OpenStudy (anonymous):

Thanks I got that from what you wrote out.

OpenStudy (wolf1728):

I just figured it might be worth mentioning. Anyway, glad to help out.